Erica invests $2744 in a savings account for 4 years that earns 1.0% compounded semi-annually.


How much money in total will she have earned at the end of 4 years?


1
Expert's answer
2021-04-20T16:24:40-0400

"A=P(1+r\/100)^n"

"r=1/2=0.5\%" "n=2\times 4=8"

"A=2744(1+0.5/100)^8=2744\times 1.0407=\$2855.6808"